OpenStudy (anonymous):

Help? Which of the following is the solution to the equation 8(y – 2) = 2y ? y = 1 y = 3 y = –3 y = –1

OpenStudy (luigi0210):

Distribute, separate and then solve.

OpenStudy (anonymous):

so far I got, 8yx16x2y

OpenStudy (luigi0210):

How did you get that?

OpenStudy (anonymous):

by doing 8 times y and 8 times 2

OpenStudy (anonymous):

hey @dreamer05 its 8y-16=2y so,6y=16 y=8/3

OpenStudy (anonymous):

how did you get 8/3?

OpenStudy (luigi0210):

@quantun No giving away answers or doing others work for them

OpenStudy (anonymous):

@dreamer05 since 6y=16 y=16/6.Divide by 2 both in numerator and denominator you will get 8/3

OpenStudy (anonymous):

but there is no option for that, unless it can be reduced? right.

OpenStudy (anonymous):

i think you should check your question once again..:P

OpenStudy (anonymous):

?

OpenStudy (luigi0210):

@dreamer05 you're right

OpenStudy (anonymous):

yeah, i mean unless its -1 but I know that isn't right.. I am a little bit confused..

OpenStudy (luigi0210):

Reduce the fraction 16/6 by 2 This is why you guide not give answers

OpenStudy (anonymous):

8/3

OpenStudy (luigi0210):

Is that one of your answer choices?

OpenStudy (anonymous):

No,.. unless its 3. but that wouldn't make since.. would it?

OpenStudy (anonymous):

@Luigi0210 what''s wrong in guiding ?

OpenStudy (luigi0210):

@quantun Nothing, I just don't think it's right for people do others work for them. And Yea qua was right, did you write that question right?

OpenStudy (anonymous):

I thought so, hold on.. 8^(y-2)=2y

OpenStudy (anonymous):

my bad..

OpenStudy (luigi0210):

That's so much better xD

OpenStudy (anonymous):

SORRRY.

OpenStudy (anonymous):

is it 8^(y-2)=2^y??

OpenStudy (anonymous):

u have posted it wrong again...

OpenStudy (anonymous):

sorry

OpenStudy (luigi0210):

Go easy on her :3

OpenStudy (anonymous):

can you solve now?

OpenStudy (luigi0210):

@dreamer05 Do you know how to start?

OpenStudy (anonymous):

Not really. I tried.

OpenStudy (luigi0210):

Well simple, get like bases.

OpenStudy (anonymous):

8 and 2?

OpenStudy (anonymous):

see 8 can be written as 2^3

OpenStudy (luigi0210):

Change 8 so it has a base of 2

OpenStudy (anonymous):

compare both sides

OpenStudy (anonymous):

what does that mean?

OpenStudy (luigi0210):

Yup, so now you have: \[2^{3(y-2)}=2^y\] Since the bases are the same, you can get rid of them and just solve: 3(y-2)=y

OpenStudy (anonymous):

so it be, 3y-6=3y

OpenStudy (anonymous):

R.H.S. will be 2y

OpenStudy (luigi0210):

Where did that 3y come from on the right?

OpenStudy (anonymous):

sorry y

OpenStudy (luigi0210):

^

OpenStudy (anonymous):

or not, idk i thought you were so suppose to add the 3 to the y. so its just 3y-6=y

OpenStudy (luigi0210):

Now just solve for y..

OpenStudy (anonymous):

-3

OpenStudy (luigi0210):

Try again

OpenStudy (anonymous):

so, just 3?

OpenStudy (anonymous):

bring y to one side and 6 to other and sovle

OpenStudy (luigi0210):

Yup

OpenStudy (anonymous):

2y=6

OpenStudy (luigi0210):

This took a bit longer than expect but hey, you got it.

OpenStudy (anonymous):

oh I did.. its 3?

OpenStudy (luigi0210):

Yup

OpenStudy (anonymous):

YAY. sorry i'm really really really bad at math...

OpenStudy (luigi0210):

No it's fine, I've dealt with you before missy ;P

OpenStudy (anonymous):

I know lolol...

OpenStudy (luigi0210):

Just keep trying and you'll be good :)

OpenStudy (anonymous):

try to solve as many problems as you can

OpenStudy (anonymous):

Thank you, thank you.

OpenStudy (luigi0210):

Welcome, welcome